A reaction mixture contains 0.15 M CS2, 0.75 M H2, 0.55 M CH4, and 0.95 M H2S.
Is this reaction at equilibrium? If it is not at equilibrium, indicate the direction in which
it much shifts to reach equilibrium. Show work to support your answer. (Hint: Do not use
Answer #1

By the given first of all we calculate the equilibrium constant

CS2(g) + 4 H2(g) \rightarrow CH4(g) + 2 H2S(g)

Equilibrium constant , K=\frac{\left [ CH_{4} \right ]\left [ H_{2}S \right ]^{2}}{\left [ CS_{2} \right ]\left [ H_{2} \right ]^{4}}

                               K=\frac{\left0.55\left \times 0.95^{2}}{\left 0.15\times \left 0.75^{4}}

                                K = 10.46

But given that K = 0.280 , the reaction is not at eqilibrium.

Since the value(10.46) obtained by our calculation is more than the actual K value ( 0.280) ,the equilibrium is shifted towards left side that is towards to the reactant side.